Explain in detail, with the aid of a graph, the two components of the demand for money.
Identify the main determinant of each component.
In your graph, illustrate each component as well as the total demand for money.
1
Expert's answer
2020-08-26T12:59:47-0400

The demand for money is the total amount of money that the population of an economy wants to hold.

 There are two components for demand for Money, they include

i) Transaction motive

ii) Precautionary Motive

The transaction motive of holding money is to take care of commercial transactions arising in the day-to-day process of living.The transaction motive for money is determined by the level of income. This can be presented as;


"M^T=kY\\
Where\ M^T=\ Transaction\ demand \ for\ money\\
Y=National\ income\\
k=constant \ of\ proportionality,\ 0<k<1"

Precautionary Motive refers to the tendency of a firm or people to hold cash, to meet the contingencies or unforeseen circumstances arising in the course of business and life, i.e sickness, a fire, unexpected business opportunity etc-all these things require a cushion of cash balances to fall back on.This type of demand for money is also determined by income and the general level of business activity.Since the precautionary demand is income-elastic, it is expressed as "M^P=f(Y)", where "M^P" is the precautionary demand for money and "f(Y)" denotes it to be the function of income.

Finally, we will combine the precautionary and transactions demand for money into a single demand schedule called the demand for active balances, which shows the total demand for money. Letting "M_A"

stand for active balances, we can therefore say


"M_A=M^P+M^T\\
Substituting\ eq.\ (1)\ into\ eq.\ (2),\ we\ then\ have\\
M_A=M^P+kY"

A plot of equation. (3) is a straight line; that is, the "M_A" demand for money is equal to constant "M^P" demand for money is equal to a constant "M^P" plus a coefficient "k" times a variable "Y" and such an equation describes a straight line, as shown in the graph below.